OpenStudy (anonymous):

How do you do this -- for algebra 2 -- question x/16-x = 2/x+2

OpenStudy (phi):

Is the problem \[ \frac{x}{16-x} = \frac{2}{x+2} \] If so, you should use parentheses: x/(16-x) = 2/(x+2) To solve, multiply both sides by (16-x) then multiply both sides by (x+2) Can you finish?